OpenStudy (anonymous):

Can someone help me please, I have to solve this equation using, foil, Fundamental Theorem of Algebra, the Rational Root Theorem, Descartes' Rule of Signs, and the Factor Theorem. 495=(x-2)(x-6)x

OpenStudy (dumbcow):

well first multiply it all out and set it equal to 0

OpenStudy (anonymous):

Ok I got 0=(x^2-8x+12)x

OpenStudy (dumbcow):

where did the 495 go? also distribute the x

OpenStudy (anonymous):

x^3-8x^2+12x=495

OpenStudy (dumbcow):

make the right side 0 by subtracting by 495

OpenStudy (anonymous):

ok so I'd get x^3-8x^2+12x=0

OpenStudy (dumbcow):

x^3 -8x^2 +12x -495 = 0

OpenStudy (anonymous):

oh ok

OpenStudy (dumbcow):

then use Descartes Thm to determine how many pos and neg roots or graph it on a calculator to guess any rational roots :)

OpenStudy (anonymous):

I don't know how to do that :/

OpenStudy (anonymous):

Thank you!

OpenStudy (dumbcow):

this shows there is 1 positive rational root looks like its 11 verify by trying to see if (x-11) factors out use synthetic division 11 | 1 -8 12 -495 _______________________ 1 you do the rest and see if remainder is 0

OpenStudy (anonymous):

yes it is

OpenStudy (dumbcow):

ok good , but the fundamental thm of algebra says we must have 3 roots because its a "x^3" function to find the other 2 roots, use quadratic formula

OpenStudy (anonymous):

ok I got +1 and -1

OpenStudy (dumbcow):

no please show work, what quadratic equation did you use?

OpenStudy (anonymous):

http://www.purplemath.com/modules/quads/qform01.gif

OpenStudy (dumbcow):

you should have this: 11 | 1 -8 12 -495 11 33 495 ------------------ 1 3 45 0 --> (x-11)(x^2 +3x +45) = 0

OpenStudy (dumbcow):

sorry the link doesn't work

OpenStudy (anonymous):

I did have that

OpenStudy (dumbcow):

\[x = \frac{-3 \pm \sqrt{9-4(1)(45)}}{2}\]

OpenStudy (anonymous):

ok I got -28.5 -171 under the division sign then divide it by 2 and I got -85.5 then divide -85.5 by -3 and 3

OpenStudy (dumbcow):

ahh, ok you can't divide by 2 because the "-171" is inside the radical \[x = \frac{-3 \pm \sqrt{-171}}{2}\]

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

whenever you have a neg num inside sqrt.....you get an imaginary number \[x = \frac{-3 \pm i \sqrt{171}}{2}\] simplify radical .... 171 = 19*9 \[x = \frac{-3 \pm 3\sqrt{19} i}{2}\]

OpenStudy (anonymous):

ok, so would that be the answer?

OpenStudy (dumbcow):

yes those are the 2 complex roots and the 1 real root is 11

OpenStudy (anonymous):

Ok Thank you